Common stock value—Zero growth Personal Finance Problem Kelsey Drums, Inc., is a well-established supplier of fine percussion(Round to the nearest cent)
Answer #1

Answer: Given data. Perpetual constant Dividend = $4.06 Requuded Required Retuon 10 years ago = 8.ly. Return custent = 4.10%
Answer #2

SOLUTION :


Dividend paid for last 15 years is $4.06 per share each year and expected to be continued in future.


Kim purchased 200 shares 10 years ago.


Let price 10 years ago be $P per share.

10 years ago, rate of return was = 8.1% = 0.081


Rate of return possible when purchased = Dividend / P 

=> 0.081 = 4.06 / P 

=> P = 4.06/0.081

=> P = Price 10 years ago = 50.12 ($) per share. 



Let price today be $P’ per share.

Today, rate of return is = 4.10% = 0.041


Rate of return possible today = Dividend / P 

=> 0.041 = 4.06 / P’ 

=> P’ = 4.06/0.041

=> P’ = Price today = 99.02 ($) per share.  


Capital gain per share = P’ - P = 99.02 - 50.12 = 48.90 ($)


So,


Total capital gain to Kim on  purchase of 200 shares 10 years ago

= 200*48.90

= 9780 ($) (ANSWER) 

answered by: Tulsiram Garg